a certain aircraft can fly 1600 miles with the wind in 5 hours and travel the same distance against the wind in 8 hours. What is the spee=d of the wind and the speed of the plane in still air? ( set two equations up ex. y=2x+3 and x=2y+10 and solve using substitution method or addition method)
Answer by mananth(16949)

against wind 8 hours
with wind 5 hours

Distance = same= 1600
plane speed =x
wind speed =y
t=d/r
1600/(x-y)=8.00
8(x-y)= 1,600.00
8x-8y=1600 ....................1
1600/ ( x + y )= 5
5 ( x + y ) = 1600
5 x + 5 y = 1600 ...............2
Multiply (1) by 5
Multiply (2) by 8
we get
40 x + -40 y = 8000
40 x + 40 y = 12800
80 x = 20800
/80
x=260 mph plane speed

plug value of x in (1)
8x-8y=1600
2080-8y= 1600
-8y=1600 -2080
-8y=-480
y=60mph wind speed